# Near-Optimal Sample Complexities of Divergence-based S-rectangular Distributionally Robust Reinforcement Learning

### v1

@@ 5.1 Robust Inventory Control Problems
Let $I$ denote the maximum inventory level, $B$ the maximum backlog, and $O$ the maximum order quantity per period. The state space is defined as $\mathcal{S}=\{-B,\cdots,0,\cdots,I\}$, the action space is $\mathcal{A}=\{0,\cdots,O\}$, where $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ denote the inventory and order item at the beginning of period $t$. Demand $D_{t}\in\{0,\cdots,D_{\max}\}$ is an i.i.d. sequence, with distribution $P_{D}\in\Delta^{D_{\max}+1}$.
@@
For our experiments, we set the parameters as follows: $I=10$, $B=5$, $O=5$, $p=3$, $c=2$, $h=0.2$, $b=3$, $\gamma=0.9$, and use the nominal demand distribution $P_{D}=[0.1,0.2,0.3,0.3,0.1]$, supported on ${0,1,2,3,4}$. For each $(s,a)$, we sample $n_{0}$ samples from the nominal transition kernels to generate the estimated transition probability $P_{n}$, and solve the DR-MDP problem with uncertainty size $\rho=1/6$ using the algorithm presented in [7].
Figure 1 illustrates the relationship between the sample size $n$ and the error $\varepsilon$ between the empirical and population value functions. As shown in the log-log plot, the slope is approximately $-0.5$ for both the KL and $\chi^{2}$ cases, indicating that the error decreases at a rate proportional to $1/\sqrt{n}$.
@@
Specifically, $\mathcal{S}$ denotes the set of all initial states, each associated with an action set $\mathcal{A}$. When an action $a_{i}\in\mathcal{A}$ is taken in state $s\in\mathcal{S}$, the system deterministically transitions (with probability 1) to the corresponding state $y_{1,s,a}\in\mathcal{Y}_{1}$. From each $y_{1,s,a}$, the system either remains in the same state with nominal probability $p$, or transitions to the corresponding absorbing state $y_{2,s,a}\in\mathcal{Y}_{2}$ with nominal probability $1-p$. All states in $\mathcal{Y}_{2}$ are absorbing, meaning that once the system enters one of these states, it remains there indefinitely via a self-loop with probability 1. The reward function is defined such that a reward of 1 is obtained only when the system is in any state within $\mathcal{Y}_{1}$; all other states yield a reward of 0. We solve the DR-MDP problem with uncertainty size $\rho=0.1$.
@@
(a) Uncertainty sets based on KL-divergence
(b) Uncertainty sets based on $\chi^{2}$-divergence
Figure 4: Estimation error versus the number of states $|\mathcal{A}|$ for the MDP instances based on the lower bound construction in Yang et al. [30].
@@ 6 Conclusion and Future Work
In this paper, we present near-optimal sample complexity results for divergence-based S-rectangular robust MDPs in the discounted reward setting. Our results are the first to achieve optimal dependence on $|\mathcal{S}|$, $|\mathcal{A}|$, and $\varepsilon$ simultaneously. We acknowledge, however, two limitations: the reliance on access to a generative model and the presence of a gap between our upper bound and the minimax lower bound established in Yang [29]. As part of future work, we aim to develop provable theoretical guarantees for other settings, including model-free algorithms and offline reinforcement learning.
@@ Appendix A Proofs of Value Function Error Bounds
In this section, we prove Proposition 1. We first show that both the population and the empirical S-rectangular Bellman operators $\mathcal{T}^{*}$ and $\hat{\mathbf{T}}^{*}$ are $\gamma$-contractions. This is a well-known fact, see for example [24]. We include a proof to make the paper self-contained.
@@
$|\mathcal{T}^{*}(v_{1})(s)-\mathcal{T}^{*}(v_{2})(s)|$
$\leq\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\left|f(v_{1})(s)-f(v_{2})(s)\right|$
$=\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\Bigg{|}\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})R(s,a,s^{\prime})+\gamma\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})v_{1}(s^{\prime})$
$-\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})R(s,a,s^{\prime})-\gamma\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})v_{2}(s^{\prime})\Bigg{|}$
$=\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\gamma\Bigg{|}\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})v_{1}(s^{\prime})-\sum_{a,s^{\prime}}\pi(a|s)P_{s,a}(s^{\prime})v_{2}(s^{\prime})\Bigg{|}$
$\leq\gamma\|\pi_{s}\|_{\infty}\|P_{s}\|_{\infty}\|v_{1}-v_{2}\|_{\infty}$
$=\gamma\|v_{1}-v_{2}\|_{\infty}$
where $\|P_{s}\|_{\infty}=\sup_{\left\|v\right\|_{\infty}=1}\left\|P_{s}v\right\|_{\infty}$ is the induced operator norm. The above inequality holds for any $s\in\mathcal{S}$, which lead to
@@
We replace $P_{s}$ with $P_{s,n}$, notice that $\|P_{s,n}\|_{\infty}\leq 1$
@@ A.1 Proof of Proposition 1
Let $v_{0}\equiv 0$ and $v_{k+1}=\hat{\mathbf{T}}^{*}(v_{k})$. $\hat{u}$ is defined as the fix point of $\hat{\mathbf{T}}^{*}$ $\hat{v}=\hat{\mathbf{T}}^{*}(\hat{v})$
$\Delta_{k+1}$ $=v_{k+1}-v^{*}$
$=\hat{\mathbf{T}}^{*}(v_{k})-\hat{\mathbf{T}}^{*}(v^{*})+\hat{\mathbf{T}}^{*}(v^{*})-\mathcal{T}^{*}(v^{*})$
$=\left[\hat{\mathbf{T}}^{*}(v^{*}+\Delta_{k})-\hat{\mathbf{T}}^{*}(v^{*})\right]+\left[\hat{\mathbf{T}}^{*}(v^{*})-\mathcal{T}^{*}(v^{*})\right]$
@@
By Lemma 3, we have
$\left\|\mathbf{H}(\Delta_{1})-\mathbf{H}(\Delta_{2})\right\|_{\infty}=\left\|\hat{\mathbf{T}}^{*}(v^{*}+\Delta_{1})-\hat{\mathbf{T}}^{*}(v^{*}+\Delta_{2})\right\|_{\infty}\leq\gamma\|\Delta_{1}-\Delta_{2}\|_{\infty},$
@@
$\leq\gamma\left\|v^{*}\right\|_{\infty}+\left\|V\right\|_{\infty}$
@@
$\leq\gamma\left(\frac{\gamma^{k-1}}{1-\gamma}+\sum_{j=0}^{k-1}\|V\|_{\infty}\right)+\left\|V\right\|_{\infty}$
$=\frac{\gamma^{k}}{1-\gamma}+\sum_{j=0}^{k}\|V\|_{\infty}.$
@@
$\|\hat{v}-v^{*}\|_{\infty}=\lim_{k\to\infty}\|\Delta_{k}\|_{\infty}\leq\sum_{j=0}^{\infty}\gamma^{j}\left\|V\right\|_{\infty}=\frac{1}{1-\gamma}\left\|\hat{\mathbf{T}}^{*}(v^{*})-\mathcal{T}^{*}(v^{*})\right\|.$
@@ B.1 Proof of Lemma 1
We first note that for each action $a$, the term $\lambda\mathbb{E}_{\overline{P}_{s,a}}[\cdot]$ is a positive constant with respect to $\eta_{a}$, while the term $-\lambda\mathbb{E}_{\overline{P}_{s,a}}[\cdot]\exp((\eta_{a}-\lambda)/\lambda)$ is concave in $\eta_{a}$, since for any $c>0$, the function $-c\exp(x)$ is concave. Moreover, the term $\sum_{a}\eta_{a}$ is affine, and hence concave. As the sum of concave functions is concave, we conclude that (B.1) is concave in $\boldsymbol{\eta}$. Next, we optimize with respect to $\boldsymbol{\eta}$ by setting the gradient with respect to each $\eta_{a}$ to zero:
$-\exp\left(\frac{\eta_{a}-\lambda}{\lambda}\right)\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(\frac{-\pi(a|s)\left(R(s,a,S)+\gamma v(S)\right)}{\lambda}\right)\right]+1=0$
@@
$\sup_{\lambda\geq 0,\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}-\lambda\sum_{a\in\mathcal{A}}\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(\frac{-\pi(a|s)\left(R(s,a,S)+\gamma v(S)\right)}{\lambda}\right)\right]-\lambda|\mathcal{A}|\rho.$
@@ B.2 Proof of Lemma 2
Since $k-1>0$ and $\lambda>0$ are constants with respect to the random variable $S$, we can factor them out of the expectation and the positive-part operator $(\cdot)_{+}$.
@@
Since $-\lambda^{-\alpha}$ is concave in $\lambda$ for any $\alpha>0$, and $\lambda|\mathcal{A}|\left(\rho+\frac{1}{k(k-1)}\right)$ is an affine function of $\lambda$, it follows that (B.3) is concave with respect to $\lambda$. To optimize over $\lambda$, we take the derivative with respect to $\lambda$ and set it to zero, which yields:
@@
Multiply $k(k-1)$ on both side of the equation, we have

### v2

@@ 5.1 Robust Inventory Control Problems
Let $I$ denote the maximum inventory level, $B$ the maximum backlog, and $O$ the maximum order quantity per period. The state space is defined as $\mathcal{S}=\{-B,\cdots,0,\cdots,I\}$, the action space is $\mathcal{A}=\{0,\cdots,O\}$, where $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ denote the inventory and order item at the beginning of period $t$. Demand $D_{t}\in\{0,\cdots,D_{\max}\}$ is an i.i.d. sequence, with distribution $P_{D}\in\Delta(D_{\max}+1)$.
@@
For our experiments, we set the parameters as follows: $I=10$, $B=5$, $O=5$, $p=3$, $c=2$, $h=0.2$, $b=3$, $\gamma=0.9$, and use the nominal demand distribution $P_{D}=[0.1,0.2,0.3,0.3,0.1]$, supported on 0,1,2,3,4{0,1,2,3,4}. For each $(s,a)$, we sample $n_{0}$ samples from the nominal transition kernels to generate the estimated transition probability $P_{n}$, and solve the DR-MDP problem with uncertainty size $\rho=1/6$ using the algorithm presented in [8].
Figure 1 illustrates the relationship between the sample size $n$ and the error $\varepsilon$ between the empirical and population value functions. As shown in the log-log plot, the slope is approximately 0.5-0.5 for both the KL and $\chi^{2}$ cases, indicating that the error decreases at a rate proportional to $1/\sqrt{n}$.
@@
$\mathcal{Y}_{1}$$\mathcal{Y}_{2}$$x_{1}$$x_{2}$$x_{|\mathcal{S}|}$$a_{1}$$a_{2}$$a_{|\mathcal{A}|}$$a_{1}$$a_{|\mathcal{A}|}$$a_{1}$$a_{|\mathcal{A}|}$1p1-$p1$-$p1$-$p1$-$p1$-$p1$-$p1$-p11111111111111pppppppppppppp
@@
Specifically, $\mathcal{S}$ denotes the set of all initial states, each associated with an action set $\mathcal{A}$. When an action $a_{i}\in\mathcal{A}$ is taken in state $s\in\mathcal{S}$, the system deterministically transitions (with probability 1) to the corresponding state $y_{1,s,a}\in\mathcal{Y}_{1}$. From each $y_{1,s,a}$, the system either remains in the same state with nominal probability $p$, or transitions to the corresponding absorbing state $y_{2,s,a}\in\mathcal{Y}_{2}$ with nominal probability 1p1-p. All states in $\mathcal{Y}_{2}$ are absorbing, meaning that once the system enters one of these states, it remains there indefinitely via a self-loop with probability 1. The reward function is defined such that a reward of 1 is obtained only when the system is in any state within $\mathcal{Y}_{1}$; all other states yield a reward of 0. We solve the DR-MDP problem with $\gamma=0.9$ and uncertainty size $\rho=0.1$.
(a) Uncertainty sets based on KL-divergence
(b) Uncertainty sets based on $\chi^{2}$-divergence
@@
Figure 4: Estimation error versus the number of actions $|\mathcal{A}|$ for the MDP instances based on the lower bound construction in Yang et al. [31].
@@ 6 Conclusion and Future Work
In this paper, we establish near-optimal sample complexity results for divergence-based S-rectangular robust MDPs under the discounted reward criterion. To the best of our knowledge, our results are the first to achieve optimal dependence on $|\mathcal{S}|$, $|\mathcal{A}|$, and $\varepsilon$ simultaneously. Looking ahead, several promising research directions remain open. A natural extension is to move beyond the generative model assumption and develop provable guarantees in more practical settings, such as model-free reinforcement learning and offline reinforcement learning. Another important avenue is to explore richer formulations of uncertainty, for instance by modeling uncertainty around the underlying noise process in distributionally robust stochastic control. Finally, it is also interesting to investigate S-rectangular robust MDPs under the average-reward criterion.
@@ Appendix A Proofs of Value Function Error Bounds
In this section, we prove Proposition 1, Corollary 1 and Corollary 2. We first show that both the population and the empirical S-rectangular Bellman operators $\mathcal{T}^{*}$ and $\hat{\mathbf{T}}^{*}$ are $\gamma$-contractions. Furthermore, for given $\pi$, the Bellman operators $\mathcal{T}^{\pi}$ and $\hat{\mathbf{T}}^{\pi}$ are also $\gamma$-contractions. This is a well-known fact, see for example [26]. We include a proof to make the paper self-contained.
Lemma 3.
Let
Then,
$\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\left|f(v_{1})(s)-f(v_{2})(s)\right|\leq\gamma\|v_{1}-v_{2}\|_{\infty}$
Proof.
Fix $s$. The reward terms cancel in the difference, hence for any $(\pi,P)$,
$|f(v_{1})(s)-f(v_{2})(s)|$ $=\gamma\Big|\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\big(v_{1}(s^{\prime})-v_{2}(s^{\prime})\big)\Big|$
$\leq\gamma\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\,\|v_{1}-v_{2}\|_{\infty}$
$=\gamma\,\|v_{1}-v_{2}\|_{\infty},$
where we used nonnegativity and normalization of $\pi(\cdot|s)$ and $P_{s,a}(\cdot)$ so the weights sum to $1$. Taking $\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}$ does not increase the right-hand side, yielding the claim. ∎
@@
$|\mathcal{T}^{*}(v_{1})(s)-\mathcal{T}^{*}(v_{2})(s)|\leq\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\left|f(v_{1})(s)-f(v_{2})(s)\right|\leq\gamma\|v_{1}-v_{2}\|_{\infty}$
The above inequality holds for any $s\in\mathcal{S}$, which lead to
@@
We replace $P_{s}$ with $P_{s,n}$, therefore
@@
Lemma 5.
$\mathcal{T}^{\pi}$ and $\hat{\mathbf{T}}^{\pi}$ are $\gamma$-contraction operators on $(\mathcal{S}\rightarrow\mathbb{R},\left\|\cdot\right\|_{\infty})$; i.e. for all $v_{1},v_{2}:\mathcal{S}\rightarrow\mathbb{R},$
$\|\mathcal{T}^{\pi}(v_{1})-\mathcal{T}^{\pi}(v_{2})\|_{\infty}\leq\gamma\|v_{1}-v_{2}\|_{\infty},$
$\|\hat{\mathbf{T}}^{\pi}(v_{1})-\hat{\mathbf{T}}^{\pi}(v_{2})\|_{\infty}\leq\gamma\|v_{1}-v_{2}\|_{\infty}.$
Proof.
By definition, we have
$|\mathcal{T}^{\pi}(v_{1})(s)-\mathcal{T}^{\pi}(v_{2})(s)|=\left|\inf_{P\in\mathcal{P}}f(v_{1})(s)-\inf_{P\in\mathcal{P}}f(v_{2})(s)\right|$
Since $|\inf_{X}f-\inf_{X}g|\leq\sup_{X}|f-g|$, we have
$|\mathcal{T}^{\pi}(v_{1})(s)-\mathcal{T}^{\pi}(v_{2})(s)|\leq\sup_{P\in\mathcal{P}}\left|f(v_{1})(s)-f(v_{2})(s)\right|\leq\sup_{\pi\in\Pi}\sup_{P\in\mathcal{P}}\left|f(v_{1})(s)-f(v_{2})(s)\right|\leq\gamma\|v_{1}-v_{2}\|_{\infty}$
The above inequality holds for any $s\in\mathcal{S}$, which lead to
$\|\mathcal{T}^{\pi}(v_{1})-\mathcal{T}^{\pi}(v_{2})\|_{\infty}\leq\gamma\|v_{1}-v_{2}\|_{\infty}.$
We replace $P_{s}$ with $P_{s,n}$, therefore,
$|\hat{\mathbf{T}}^{\pi}(v_{1})(s)-\hat{\mathbf{T}}^{\pi}(v_{2})(s)|$ $\leq\gamma\|\pi_{s}\|_{\infty}\|P_{s,n}\|_{\infty}\|v_{1}-v_{2}\|_{\infty}$
$=\gamma\|v_{1}-v_{2}\|_{\infty}.$
which lead to
$\|\hat{\mathbf{T}}^{\pi}(v_{1})-\hat{\mathbf{T}}^{\pi}(v_{2})\|_{\infty}\leq\gamma\|v_{1}-v_{2}\|_{\infty}.$
@@ A.1 Proof of Proposition 1
Let $u_{0}\equiv 0$ and $_{k+1}=\hat{T}(u_{k}). u^\hat{u}$ and $u^{*}$ is defined as the fix point of $\hat{T}$ and $T$ respectively.
$\Delta_{k+1}$ $=u_{k+1}-u^{*}$
$=\hat{T}(u_{k})-\hat{T}(u^{*})+\hat{T}(u^{*})-T(u^{*})$
$=\left[\hat{T}(u^{*}+\Delta_{k})-\hat{T}(u^{*})\right]+\left[\hat{T}(u^{*})-T(u^{*})\right]$
@@
Since $\hat{T}$ is a $\gamma$-contraction operator, we have
$\left\|\mathbf{H}(\Delta_{1})-\mathbf{H}(\Delta_{2})\right\|_{\infty}=\left\|\hat{T}(u^{*}+\Delta_{1})-\hat{T}(u^{*}+\Delta_{2})\right\|_{\infty}\leq\gamma\|\Delta_{1}-\Delta_{2}\|_{\infty},$
@@
$\leq\gamma\left\|u^{*}\right\|_{\infty}+\left\|V\right\|_{\infty}$
@@
$\leq\gamma\left(\frac{\gamma^{k-1}}{1-\gamma}+\sum_{j=0}^{k-1}\gamma^{j}\|V\|_{\infty}\right)+\left\|V\right\|_{\infty}$
$=\frac{\gamma^{k}}{1-\gamma}+\sum_{j=0}^{k}\gamma^{j}\|V\|_{\infty}.$
@@
$\|\hat{u}-u^{*}\|_{\infty}=\lim_{k\to\infty}\|\Delta_{k}\|_{\infty}\leq\sum_{j=0}^{\infty}\gamma^{j}\left\|V\right\|_{\infty}=\frac{1}{1-\gamma}\left\|\hat{T}(u^{*})-T(u^{*})\right\|.$
A.2 Proof of Corollary 1
Proof.
By Lemma 4, both $\mathcal{T}^{*}$ and $\hat{\mathbf{T}}^{*}$ are $\gamma$-contractions. Applying Proposition 1 with $\mathcal{T}^{*}, T^=𝐓^\hat$$\hat{\mathbf{T}}^{*}$ and $u^{*}=v^{*}$ yields
$\|\hat{v}-v^{*}\|_{\infty}\;\leq\;\frac{1}{1-\gamma}\,\|\hat{\mathbf{T}}^{*}(v^{*})-\mathcal{T}^{*}(v^{*})\|_{\infty},$
which proves the statement. ∎
A.3 Proof of Corollary 2
Proof.
[…]
$\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\|_{\infty}\;\leq\;\frac{1}{1-\gamma}\,\big\|\hat{\mathbf{T}}^{\pi}\!\left(V_{\mathcal{P}}^{\pi}\right)-\mathcal{T}^{\pi}\!\left(V_{\mathcal{P}}^{\pi}\right)\big\|_{\infty}.$
Since the above inequality holds for any $\pi\in\Pi$, taking the supremum over $\pi\in\Pi$ on both sides yields
$\sup_{\pi\in\Pi}\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\|_{\infty}\;\leq\;\frac{1}{1-\gamma}\,\sup_{\pi\in\Pi}\big\|\hat{\mathbf{T}}^{\pi}\!\left(V_{\mathcal{P}}^{\pi}\right)-\mathcal{T}^{\pi}\!\left(V_{\mathcal{P}}^{\pi}\right)\big\|_{\infty},$
which proves the claim. ∎
@@ B.1 Proof of Lemma 1
We first note that for each action $a$, the term $\lambda\mathbb{E}_{\overline{P}_{s,a}}[\cdot]$ is a positive constant with respect to $\eta_{a}$, while the term $\lambda\mathbb{E}_{\overline{P}_{s,a}}[\cdot]\exp((\eta_{a}-\lambda)/\lambda)$ is concave in $\eta_{a}$, since for any $c>0$, the function $c\exp(x)$ is concave. Moreover, the term $\sum_{a}\eta_{a}$ is affine, and hence concave. As the sum of concave functions is concave, we conclude that (B.1) is concave in $\boldsymbol{\eta}$. Next, we optimize with respect to $\boldsymbol{\eta}$ by setting the gradient with respect to each $\eta_{a}$ to zero:
$\exp\left(\frac{\eta_{a}-\lambda}{\lambda}\right)\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(\frac{-\pi(a|s)\left(R(s,a,S)+\gamma v(S)\right)}{\lambda}\right)\right]+1=0$
@@
$\sup_{\lambda\geq 0,\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}-\lambda\sum_{a\in\mathcal{A}}\log\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(\frac{-\pi(a|s)\left(R(s,a,S)+\gamma v(S)\right)}{\lambda}\right)\right]-\lambda|\mathcal{A}|\rho.$
@@ B.2 Proof of Lemma 2
Since k1>0k-1>0 and $\lambda>0$ are constants with respect to the random variable $S$, we can factor them out of the expectation and the positive-part operator $(\cdot)_{+}$.
@@
Since $\lambda^{-\alpha}$ is concave in $\lambda$ for any $\alpha>0$, and $\lambda|\mathcal{A}|\left(\rho+\frac{1}{k(k-1)}\right)$ is an affine function of $\lambda$, it follows that (B.3) is concave with respect to $\lambda$. To optimize over $\lambda$, we take the derivative with respect to $\lambda$ and set it to zero, which yields:
@@
Multiply k(k1)k(k-1) on both side of the equation, we have